# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ Abstract
Let $\mathbb{D}$ be a division ring with finite degree $d$ over a central subfield $\mathbb{F}$. Under mild cardinality assumptions on $\mathbb{F}$, we determine the greatest possible dimension for an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ in which no matrix has a nonzero fixed vector, and we elucidate the structure of the spaces that have the greatest dimension. The classification involves the associative composition algebras over $\mathbb{F}$ and requires the study of intransitive spaces of operators between finite-dimensional vector spaces.
These results are applied to solve various problems of the same flavor, including the determination of the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ in which all the matrices have rank at least $r$ (for some fixed $r$), as well as the structure of the spaces that have the greatest possible dimension if $n=p=r$; and finally the structure of large $\mathbb{F}$-linear subspaces of $\mathbb{F}$-diagonalisable matrices is elucidated.
@@
Keywords: affine space, rank, matrices over division rings, intransitivity, trivial spectrum spaces, simultaneous reduction, associative composition algebras
List of notation
[…]
@@ 1.1 The problem
Throughout, we let $\mathbb{D}$ be a (unital and associative) division ring, $\mathbb{F}$ be a subfield of the center of $\mathbb{D}$, and we assume that $\mathbb{D}$ has finite degree $d\geq 1$ over $\mathbb{F}$. All the vector spaces over $\mathbb{D}$ we manipulate will be finite-dimensional and right vector spaces, with the notable exception of dual spaces which are naturally left vector spaces. This way, we can safely practice matrix computations in the usual sense, with vectors represented by columns.
The space $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ of all endomorphisms of the $\mathbb{D}$-vector space $V$ has no natural structure of $\mathbb{D}$-vector space (because we are avoiding any commutativity assumption), but it can nevertheless be seen as a vector space over $\mathbb{F}$. An element $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is said to have $\mathbb{F}$-trivial spectrum when it has no nonzero (right) eigenvalue in $\mathbb{F}$ (we adopt a similar definition for matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$). A subset of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is said to have $\mathbb{F}$-trivial spectrum whenever all its elements have $\mathbb{F}$-trivial spectrum. For an $\mathbb{F}$-linear subspace $S$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, to have $\mathbb{F}$-trivial spectrum is equivalent to the condition that no element of $S$ has $1$ as eigenvalue, i.e., no element of $S$ has a nonzero fixed vector.
In matrix terms, an example of large $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with $\mathbb{F}$-trivial spectrum is the set of all upper-triangular matrices
$\begin{bmatrix}x_{1}&&*\\ &\ddots&\\ (0)&&x_{n}\end{bmatrix}$
in which the $i$-th diagonal entry $x_{i}$ is required to belong to a given $\mathbb{F}$-linear hyperplane $H_{i}$ of $\mathbb{D}$ such that $\mathbb{F}\oplus H_{i}=\mathbb{D}$. In case $\mathbb{D}=\mathbb{F}$, all the hyperplanes $H_{i}$ are zero and we have simply described the strictly upper-triangular matrices over $\mathbb{D}$. Note that the dimension over $\mathbb{F}$ of the above space is
$\alpha(n,d):=n(d-1)+d\,\frac{n(n-1)}{2},$
and it will be regularly useful to observe that
$\forall m,\;\forall n,\quad\alpha(m+n,d)=\alpha(m,d)+\alpha(n,d)+mnd.$
Trivial spectrum subspaces have been studied in the last decade [15, 18], but only in the special case $\mathbb{D}=\mathbb{F}$. One motivation for considering such spaces is that they are a natural variation of linear subspaces of nilpotent operators: it is unavoidable to cite the celebrated Gerstenhaber theorem [7], which states that the greatest possible dimension for a linear subspace of nilpotent $n$-by-$n$ matrices over a field is $\frac{n(n-1)}{2}$, with equality attained only for the conjugates of the subspace of all strictly upper-triangular matrices. See [17] for a generalization to arbitrary division rings, and [9, 26] for alternative proofs.
There is a more profound justification for studying trivial spectrum subspaces: they are intimately connected with $\mathbb{F}$-affine subspaces of invertible elements of the ring $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Let indeed $\mathcal{S}$ be an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, with translation vector space denoted by $\overrightarrow{S}$ (hence $\overrightarrow{S}$ is a $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$). Assume that $\mathcal{S}$ contains $\operatorname{\operatorname{id}}_{V}$. Then $\mathcal{S}\subseteq\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$ if and only if $\operatorname{\operatorname{id}}_{V}-u$ is invertible for all $u\in\overrightarrow{S}$, i.e., no element of $\overrightarrow{S}$ has a nonzero fixed vector. As noted in the above, this is equivalent to the condition that $\overrightarrow{S}$ has $\mathbb{F}$-trivial spectrum. It follows that the greatest possible dimension for an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is exactly the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that is included in $\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$. And by a method due to Meshulam [10], this can also be used to obtain the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ in which all the elements have their rank greater than or equal to some fixed $r\in\mathopen{[\![}1,\min(n,p)\mathclose{]\!]}$.
In [23], we have also discovered a surprising connection between $\mathbb{F}$-trivial spectrum subspaces (when $\mathbb{D}=\mathbb{F}$) and the apparently opposite notion of an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, meaning an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ in which every element is diagonalisable over $\mathbb{F}$ (i.e., every element is represented in some basis by a diagonal matrix with all diagonal entries in $\mathbb{F}$, or alternatively is annihilated by a polynomial of $\mathbb{F}[t]$ that is split with simple roots over $\mathbb{F}$). In particular, it was shown that the classification of the trivial spectrum subspaces over $\mathbb{R}$ (there, $\mathbb{D}=\mathbb{F}=\mathbb{R}$) with the greatest possible dimension yields the one of the $\mathbb{R}$-diagonalisable linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ with maximal dimension, which simply turn out to be the conjugates of the subspace $\operatorname{\operatorname{S}}_{n}(\mathbb{R})$ of all $n$-by-$n$ symmetric matrices. As a special case of the results on $\mathbb{F}$-trivial spectrum subspaces, we will obtain both the Hermitian and quaternionic cases, and in particular we will prove that every $\mathbb{R}$-diagonalisable linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ has dimension over $\mathbb{R}$ at most $n^{2}$, with equality attained only for the conjugates of the subspace $\mathcal{H}_{n}(\mathbb{C})$ of all $n$-by-$n$ (complex) Hermitian matrices.
1.2 Reduction to the irreducible case
An $\mathbb{F}$-trivial spectrum (linear) subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (or of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$) is called optimal when it has the greatest possible dimension among such spaces. A subset $\mathcal{X}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is called irreducible when $V$ is nonzero and $\mathcal{X}$ has no nontrivial invariant $\mathbb{D}$-linear subspace, and is called reducible otherwise. The study of optimal $\mathbb{F}$-trivial spectrum subspaces can entirely be reduced to the one of irreducible optimal subspaces. The reduction process is not specific to $\mathbb{F}$-trivial spectrum subspaces of endomorphisms, and is actually shared by many properties on matrices: an account of this is given in the appendix of [25], and here we will simply retrieve the results. Key to this is the observation that the $\mathbb{F}$-trivial spectrum property satisfies the following two conditions:
  • It is invariant under similarity, i.e., if $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ has this property then so does $\varphi\circ u\circ\varphi^{-1}$ for every isomorphism $\varphi:V\overset{\simeq}{\longrightarrow}V^{\prime}$ of $\mathbb{D}$-vector spaces;
  • It satisfies the strong inheritance condition, meaning that for every $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ and every nontrivial $\mathbb{D}$-linear subspace $W$ of $V$ that is invariant under $u$, the endomorphism $u$ has $\mathbb{F}$-trivial spectrum if and only if the endomorphisms of $W$ and $V/W$ induced by $u$ have $\mathbb{F}$-trivial spectrum (this follows from the observation that $u$ can be seen as an endomorphism of the $\mathbb{F}$-vector space $V$, which is finite-dimensional).
By lemma A.1 in [25], we deduce the following result:
Proposition 1.1.
Let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Then the $\mathcal{S}$-invariant subspaces of $V$ are totally ordered under inclusion.
A flag of ($\mathbb{D}$-linear subspaces of) $V$ is an increasing list $V_{0}\subset\cdots\subset V_{p}$ of $\mathbb{D}$-linear subspaces with $V_{0}=\{0\}$ and $V_{p}=V$. Beware that we do not require that $\dim_{\mathbb{D}}V_{i}=i$ for all $i\in\mathopen{[\![}0,p\mathclose{]\!]}$ (in which case we would say that this flag is complete). The previous result can then be restated as saying that when $\mathcal{S}$ is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, the $\mathcal{S}$-invariant subspaces form a flag.
Next, let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, and consider the $\mathcal{S}$-invariant subspaces $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$. Given respective subsets $\mathcal{X}_{1},\dots,\mathcal{X}_{p}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{1}/V_{0}),\dots,\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{p}/V_{p-1})$, one denotes by $\mathcal{X}_{1}\vee\cdots\vee\mathcal{X}_{p}$ the set of all $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that leave all the $V_{i}$’s invariant and induce, for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, an element of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$ that belongs to $\mathcal{X}_{i}$. Then, as a consequence of theorem A.2 in [25], we obtain:

### v2

# Affine subspaces of units in simple algebras

@@ Abstract
Let $\mathcal{A}$ be a simple algebra over a field $\mathbb{F}$. Under a mild cardinality assumption on $\mathbb{F}$, we determine the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\mathcal{A}$ that is included in the group of units $\mathcal{A}^{\times}$, and we describe the spaces that have the greatest possible dimension. This is equivalent to the problem of determining the greatest possible dimension for an $\mathbb{F}$-linear subspace $S$ of $\mathcal{A}$ in which $x-1_{\mathcal{A}}$ is a unit for all $x\in S$, and we elucidate the structure of these linear subspaces up to conjugation when their dimension reaches the greatest possible one.
These classifications involve the associative composition algebras over $\mathbb{F}$. Over fields of characteristic other than $2$, the first problem is essentially reduced to the classification of nonisotropic quadratic forms over $\mathbb{F}$ and of nonisotropic Hermitian forms over quadratic and quaternionic extensions of $\mathbb{F}$.
These results are intimately connected with the problem of intransitive operator spaces between finite-dimensional vector spaces over division rings, which we study in depth: in particular, we generalize a dual version of Atkinson’s theorem on primitive spaces of bounded rank matrices.
@@
Keywords: simple algebras, affine spaces, matrices over division rings, intransitivity, trivial spectrum spaces, associative composition algebras
@@ 1.1 The problem
Let $\mathbb{F}$ be a field. In the seminal [5], Murray Gerstenhaber proved that every linear subspace of nilpotent $n$-by-$n$ matrices (over $\mathbb{F}$) has its dimension no larger than $\frac{n(n-1)}{2}$, and equality is attained only for the conjugates of the space $\operatorname{\operatorname{NT}}_{n}(\mathbb{F})$ of all strictly upper-triangular matrices (Gerstenhaber actually required that $|\mathbb{F}|\geq n$, but his result has been later revealed to hold over an field [20]).
[…]
Meshulam’s problem was solved in the early 2010’s. In [9] and [10], the bound $\frac{n(n-1)}{2}$ from Gerstenhaber’s theorem was generalized to trivial spectrum subspaces over any field, and then a complete classification of the trivial spectrum subspaces with dimension $\frac{n(n-1)}{2}$, called the optimal ones, was achieved [12] over any field with cardinality greater than $2$. Spectacularly, the classification is intimately connected with the quadratic structure of the field $\mathbb{F}$: the classification of the optimal trivial spectrum subspaces is reduced to the classification of nonisotropic bilinear forms over finite-dimensional $\mathbb{F}$-vector spaces, up to congruence (i.e., equivalence and multiplication with a nonzero scalar). In particular, for quadratically closed fields the trivial spectrum subspaces of maximal dimension all consist of nilpotent operators. As a consequence, all the affine subspaces of invertible elements of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ have dimension at most $\frac{n(n-1)}{2}$, and (for fields with more than $2$ elements) the classification up to equivalence of the ones of maximal dimension, called the optimal ones, has been reduced to the classification of nonisotropic quadratic forms over $\mathbb{F}$ up to congruence. To be more precise, to each equivalence class of optimal affine subspace of invertible elements of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ can be attached a list $(q_{1},\dots,q_{p})$ of nonisotropic quadratic forms on finite-dimensional vector spaces over $\mathbb{F}$ such that $n=\sum_{k=1}^{p}\dim q_{k}$, and the equivalence class of the said matrix space determines the congruence class of each $q_{i}$. We must of course insist that this result fails for fields with $2$ elements. For example, for $\mathbb{F}=\mathbb{R}$ this shows that there are as many equivalence classes of optimal affine spaces as there are ordered partitions of $n$. Finally, these results have been applied to the study of affine subspaces of matrices with rank bounded below [11] and to the one of linear subspaces of diagonalisable matrices [18].
Here we consider the more general problem of affine subspaces of units in simple algebras. A simple algebra over $\mathbb{F}$ is a nontrivial unital associative $\mathbb{F}$-algebra $\mathcal{A}$ with finite dimension over $\mathbb{F}$ and which has no nontrivial two-sided ideal. Two subsets $\mathcal{S}$ and $\mathcal{T}$ of $\mathcal{A}$ are called equivalent whenever there exist units $a,b$ in $\mathcal{A}^{\times}$ such that $\mathcal{T}=a\mathcal{S}b$, in which case we write $\mathcal{S}\sim\mathcal{T}$, and they are called similar whenever there exists a unit $a\in\mathcal{A}^{\times}$ such that $\mathcal{T}=a\mathcal{S}a^{-1}$, in which case we write $\mathcal{S}\simeq\mathcal{T}$.
By a classical theorem of Wedderburn, the simple $\mathbb{F}$-algebras are the isomorphs of the matrix algebras of the form $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, where $\mathbb{D}$ is a division ring which contains $\mathbb{F}$ as a central subfield (i.e., $\mathbb{F}$ is included in the center of $\mathbb{D}$) and has finite dimension over $\mathbb{F}$, and $n$ is a positive integer. Here, we expressly do not require that $\mathcal{A}$ is central as an $\mathbb{F}$-algebra, for this would leave out many interesting cases, e.g., the one of the $\mathbb{R}$-algebra $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$.
Our main problem is the extension of Meshulam’s: Given a simple $\mathbb{F}$-algebra $\mathcal{A}$, what is the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\mathcal{A}$ that is included in the unit group $\mathcal{A}^{\times}$? And what are the ones that have the greatest possible dimension?
We can easily translate Meshulam’s observation in this general situation: every $\mathbb{F}$-affine subspace of $\mathcal{A}$ is equivalent to one which contains the unity $1_{\mathcal{A}}$, and hence we can largely bound the discussion by studying the $\mathbb{F}$-linear subspaces $S$ of $\mathcal{A}$ for which $1_{A}-x$ is a unit for all $x\in S$.
Definition 1.1.
A subset of $\mathcal{A}$ is called unital when it is included in the group $\mathcal{A}^{\times}$ of units of $\mathcal{A}$. A unital $\mathbb{F}$-affine subspace is called optimal when it has the greatest possible dimension among such subspaces.
An $\mathbb{F}$-linear subspace $S$ of $\mathcal{A}$ is said to have trivial spectrum whenever $1_{\mathcal{A}}+S$ is unital, i.e., $1_{\mathcal{A}}-x\in\mathcal{A}^{\times}$ for all $x\in S$. It is called an optimal trivial spectrum subspace whenever it has the greatest possible dimension among such subspaces.
We note that, at the other end of the spectrum, the analogue problem for the $\mathbb{F}$-affine subspaces of elements of $\mathcal{A}$ that contain no unit has been entirely solved in [16] (the greatest possible dimension has been obtained, as well as the structure of the spaces with greatest possible dimension). It can hardly be disputed that this problem is considerably simpler than the one we are tackling here.
1.2 Reduction to the irreducible spaces
In order to state our results more clearly, it is unavoidable to use Wedderburn’s structure theorem for simple algebras and to frame the problem into one on spaces of linear operators or of matrices. So, we now let $\mathbb{D}$ be an arbitrary division ring which contains $\mathbb{F}$ as a central subfield, and with finite dimension $d$ over $\mathbb{F}$. We will consider right vector spaces over $\mathbb{D}$ throughout, enabling us to represent $\mathbb{D}$-linear mappings between vector spaces by matrices with entries in $\mathbb{D}$ in the usual way (so that vector are represented by columns and applying an operator corresponds to multiplying the coordinate vector on the left by the matrix of the operator). For such vector spaces $U$ and $V$, we get the $\mathbb{F}$-vector space $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ of all $\mathbb{D}$-linear mappings from $U$ to $V$, and the simple $\mathbb{F}$-algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(U)$ of all endomorphisms of the $\mathbb{D}$-vector space $U$. The latter is isomorphic to the matrix $\mathbb{F}$-algebra $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ where $n:=\dim_{\mathbb{D}}U$ through a choice a basis of $U$. Throughout, we consider $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ as an $\mathbb{F}$-vector space and will not consider potential enriched structures (for example, when $\mathbb{D}=\mathbb{C}$ and $\mathbb{F}=\mathbb{R}$, we only view $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ as an $\mathbb{R}$-vector space, and never as a $\mathbb{C}$-vector space).
We can immediately give an example of a very large trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, which generalizes Gerstenhaber’s nilpotent spaces: we can choose various $\mathbb{F}$-linear hyperplanes $H_{1},\dots,H_{n}$ of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$, and form the set $H_{1}\vee\cdots\vee H_{n}$ of all matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ of the form
$\begin{bmatrix}h_{1}&&[*]\\ &\ddots&\\ (0)&&h_{n}\end{bmatrix}$
where $h_{i}\in H_{i}$ for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, and the entries above the diagonal are arbitrary. It is clear that this is a trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. We note that this space has dimension
$\alpha(n,d):=n(d-1)+d\binom{n}{2}$.
More generally, we can create new trivial spectrum subspaces of matrices from known ones thanks to the junction operator $\vee$. For nonempty subsets $\mathcal{M}_{1}\subseteq\operatorname{\operatorname{M}}_{n_{1}}(\mathbb{D}),\dots,\mathcal{M}_{p}\subseteq\operatorname{\operatorname{M}}_{n_{p}}(\mathbb{D})$, we define their joint $\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}$ as the set of all block upper-triangular matrices of the form
$\begin{bmatrix}M_{1}&&[*]\\ &\ddots&\\ [0]&&M_{p}\end{bmatrix}$
where $M_{1}\in\mathcal{M}_{1},\dots,M_{p}\in\mathcal{M}_{p}$ and the blocks above the diagonal are arbitrary. Since $\mathbb{D}$ is finite-dimensional over $\mathbb{F}$, for any block upper-triangular matrix $M=\begin{bmatrix}A&B\\ [0]&C\end{bmatrix}$ with entries in $\mathbb{D}$ the invertibility of $M$ is equivalent to the one of $A$ and $C$. As a consequence, $\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}$ is a unital $\mathbb{F}$-affine subspace if and only if each $\mathcal{M}_{i}$ is a unital $\mathbb{F}$-affine subspace, and $\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}$ is a trivial spectrum linear subspace if and only if each $\mathcal{M}_{i}$ is a trivial spectrum linear subspace.
The above construction is of course intimately connected with the notion of irreducibility for subsets of endomorphisms. Given a finite-dimensional vector space $V$ over $\mathbb{D}$ and a subset $\mathcal{X}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, a $\mathbb{D}$-linear subspace $V_{0}$ of $V$ is called $\mathcal{X}$-invariant whenever $\forall f\in\mathcal{X},\;\forall x\in V_{0},\;f(x)\in V_{0}$, and we say that $\mathcal{X}$ is irreducible whenever $V\neq\{0\}$ and the only $\mathcal{X}$-invariant subspaces of $V$ are $\{0\}$ and $V$. If $S$ is a trivial spectrum subspace and $V_{0}$ is an $S$-invariant subspace of $V$, then $S$ induces respective linear subspaces of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(V_{0})$ and $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(V/V_{0})$, and both are trivial spectrum subspaces.
Then, we have the equivalent of the junction operator for subsets of endomorphisms: given a flag $(V_{0},\dots,V_{p})$ of $\mathbb{D}$-linear subspaces of $V$, and given (non-empty) subsets $\mathcal{S}_{1}\subseteq\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{1}/V_{0}),\dots,\mathcal{S}_{p}\subseteq\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{p}/V_{p-1})$, we denote by $\mathcal{S}_{1}\vee\cdots\vee\mathcal{S}_{p}$ the set of all $\mathbb{D}$-linear endomorphisms of $V$ that leave each space $V_{i}$ invariant and induce, for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, an endomorphism of $V_{i}/V_{i-1}$ that belongs to $\mathcal{S}_{i}$.
We can now invoke a result that was obtained in the appendix of [19] as a special case of a more general one: